# A Phase-field Model for Apoptotic Cell Death

### v2

@@ 4. Qualitative comparison of electron microscopy images and simulations
For a qualitative comparison between our apoptotic simulations and electron microscopy images, we used binarised images from You et al. [you2015kml001]. These show prostate cancer cells exposed to an anti-cancer drug, KML001, which is known to induce apoptosis. We compared and identified similar structures and topological transitions between the electron microscopy and the simulation images to demonstrate the model’s capabilities. The cyto phase field $\varphi$ in our model setup was compared to electron microscopy images of prostate cancer cell lines PC3, DU145, and LNCaP [you2015kml001] to determine whether similar morphological changes were taking place in both systems. In this comparison, KML001 was approximated to act as the apoptotic phase field $\sigma$; however, this phase field cannot be visually compared to KML001, so we focused on the comparison between the cyto phase field $\varphi$ and cell images.
@@
As mentioned previously, the predominant focus of our model was placed on simulating key apoptotic features: finger formation, cell shrinkage, fragmentation and nucleation. In our simulations, the cyto phase field $\varphi$ exhibits all characteristics. Starting with finger formation, we provide a binarised version of an electron microscopy image exhibiting this feature from [you2015kml001]. To supplement the electron microscopy image demonstrating this characteristic, we include a simulation from our model that provides similar features in Figure 11. As can be seen from the comparisons, finger formation is present in both the microscopy image of a PC3 cell (A) and the simulation (B), as well as the nucleation. Thus, thermodynamic processes described by our model also occur when the cancerous cell is exposed to the anti-cancer drug. (A) differs from (B) in the aspect of fragmentation, with more pronounced cell fragments around the prostate cancer cell present in the microscopy image. However, it can be seen in Figure 4 that fragments of the cyto phase field $\varphi$ can form during the degradation of the phase field as a result of apoptosis. In this simulation, the fragmentation appeared as the parameter dictating the interface width $\tilde{\epsilon}$ and thus the dimensionless parameter $\ell_{\varphi}$ was increased. Fragmentation of the phase field $\varphi$ was also observed when the reaction rate parameters $k_{1}$ was increased and $k_{2}$ was decreased, see Figures 6 and 7. Thus, depending on the experimental conditions, different cellular morphologies can be observed, hence explaining the variation in the comparisons.
Electron Microscopy Image Phase-field $\varphi(\boldsymbol{x},t)$
@@
Another key characteristic of apoptosis observed in the electron microscopy images in [you2015kml001] is the nucleation, that is, the formation of cavities. Figure 11 exhibits the nucleation in the electron microscopy image (A) and the simulation (B). However, a more pronounced representation of nucleation can be seen in Figure 12, where two large cavities can be seen at the interface in both the microscopy image of a PC3 cell (A) and the simulation (B). In both images, it is observed that the structure surrounding the nucleation is maintained over a greater time course than the integrity of the primary region of the cyto phase field $\varphi$. This could be explained in our simulations by the nonexistence of the apoptotic phase field $\sigma$. Thus, further degeneration of the cyto phase field $\varphi$ surrounding the nucleation could not occur, and the structure around the nucleation persists. Furthermore, it was observed in Figure 9 that the internal configurational force at the interface vanishes around the nucleation and only exists where the phase field $\sigma$ is still present.
Electron Microscopy Image Phase-field $\varphi(\boldsymbol{x},t)$
@@
In the electron microscopy images presented by You et al. [you2015kml001], the LNCaP cell line noticeably forms a greater number of cavities in the cell, and the cavities appear in higher frequency compared to those in PC3 cells. This can be seen in Figure 13 and is compared to the results from our simulation seen in Figure 8. Thus, we demonstrate we can account for these morphological changes in our simulations by introducing a higher rate of reaction for the degradation of the cyto phase field $\varphi$.
Electron Microscopy Image Phase-field $\varphi(\boldsymbol{x},t)$
@@ 5. Conclusions
Apoptosis proceeds through two well-established signalling routes, that is, the intrinsic, mitochondria-mediated pathway and the extrinsic, receptor-mediated pathway, as discussed by Lossi [Lossi2022_apoptosis] and Mustafa et al. [Mustafa2024_ApoptosisOverview]. Both mechanisms ultimately result in the activation of caspases, which coordinate the enzymatic disassembly of the cytoskeleton, organelles, and nucleus. In our formulation, this biochemical activation is represented by a spatially distributed apoptotic phase, coupled to a cyto phase describing the cell itself. The model treats the activation field in a pathway-agnostic manner, focusing on its mechanical and morphological consequences rather than its molecular origin.
We developed a biochemo–mechanical phase-field framework, given by the system in Equations (12), and implemented it in the Dedalus package [burns2020dedalus]. Variational derivation yielded a coupled set of reaction–diffusion and configurational-mechanics equations capable of reproducing characteristic apoptotic morphologies and topological transitions. Simulations of the cyto phase $\varphi$ interacting with the apoptotic phase $\sigma$ reveal finger formation, nucleation, fragmentation, and shrinkage. The parameters $\ell_{\varphi}$, $\ell_{r}^{\varphi}$, $\ell_{r}^{\sigma}$, $k_{2}$, and $\ell_{2}$ control the onset and extent of these topological transitions: decreasing $\tilde{\epsilon}$ suppresses fragmentation, while larger $\ell_{r}^{\varphi}$, $\ell_{r}^{\sigma}$, $k_{2}$, and $\ell_{2}$ enhance cavity and finger formation. Internal configurational-force analysis identifies zones of highest degradation, matching the protrusive regions observed experimentally.
Comparison with electron microscopy images from You et al. [you2015kml001], where KML001 induces apoptosis and autophagic cell death in prostate cancer cells via oxidative stress, shows morphological agreement to a great extent. The model captures the same qualitative sequence of surface deformation, cavity formation, and eventual fragmentation observed in PC3 and LNCaP cells, confirming its ability to reproduce experimentally reported apoptotic morphologies. The discrepancies seen between the experimental and simulated data are due to the highly heterogeneous nature of the cytoplasm and the cell as a whole. Furthermore, the simulation utilises a simple initial condition where the cell is assumed to be a circular disk, in reality, cells appear in a myriad of shapes and sizes.
@@ Appendix A Configurational mechanics
The evolution of the interfaces may be described by configurational mechanics. This approach, proposed by Gurtin [gurtin1999configurational] generalises the concept of Newtonian forces. Configurational forces are connected to the integrity of the material in a broad sense, explaining the evolution of material and immaterial interfaces. In what follows, we use the configurational balance suggested by Fried [fried2006relationship] and Espath & Calo [espath2021phase].

### v3

@@ 4. Comparison of electron microscopy images and simulations
For a qualitative comparison between our apoptotic simulations and electron microscopy images, we used binarised images from You et al. [25]. These show prostate cancer cells exposed to an anti-cancer drug, KML001, which is known to induce apoptosis. We compared and identified similar structures and topological transitions between the electron microscopy and the simulation images to demonstrate the model’s capabilities. The cyto phase field $\varphi$ in our model setup was compared to electron microscopy images of prostate cancer cell lines PC3, DU145, and LNCaP [25] to determine whether similar morphological changes were taking place in both systems. In this comparison, KML001 was approximated to act as the cytotoxic phase field $\sigma$. However, this phase field cannot be visually compared to KML001, so we focused on the comparison between the cyto phase field $\varphi$ and cell images.
@@
As mentioned previously, the predominant focus of our model was placed on simulating key apoptotic features: finger formation, cell shrinkage, fragmentation and nucleation. In our simulations, the cyto phase field $\varphi$ exhibits all characteristics. Starting with finger formation, we provide a binarised version of an electron microscopy image exhibiting this feature from [25]. To supplement the electron microscopy image demonstrating this characteristic, we include a simulation from our model that provides similar features in Figure 11. As can be seen from the comparisons, finger formation is present in both the microscopy image of a PC3 cell (A) and the simulation (B), as well as the nucleation. Thus, thermodynamic processes described by our model also occur when the cancerous cell is exposed to the anti-cancer drug. (A) differs from (B) in the aspect of fragmentation, with more pronounced cell fragments around the prostate cancer cell present in the microscopy image. However, it can be seen in Figure 4 that fragments of the cyto phase field $\varphi$ can form during the degradation of the phase field as a result of apoptosis. In this simulation, the fragmentation appeared as the parameter dictating the interface width $\tilde{\epsilon}$ and thus the dimensionless parameter $\ell_{\varphi}$ was increased. Fragmentation of the phase field $\varphi$ was also observed when the reaction rate parameters $k_{1}$ was increased and $k_{2}$ was decreased, see Figures 6 and 7. Thus, depending on the experimental conditions, different cellular morphologies can be observed, hence explaining the variation in the comparisons.
Electron Microscopy Image Phase field $\varphi(\boldsymbol{x},t)$
@@
Another key characteristic of apoptosis observed in the electron microscopy images in [25] is the nucleation, that is, the formation of cavities. Figure 11 exhibits the nucleation in the electron microscopy image (A) and the simulation (B). However, a more pronounced representation of nucleation can be seen in Figure 12, where two large cavities can be seen at the interface in both the microscopy image of a PC3 cell (A) and the simulation (B). In both images, it is observed that the structure surrounding the nucleation is maintained over a greater time course than the integrity of the primary region of the cyto phase field $\varphi$. This could be explained in our simulations by the nonexistence of the cytotoxic phase field $\sigma$. Thus, further degeneration of the cyto phase field $\varphi$ surrounding the nucleation could not occur, and the structure around the nucleation persists. Furthermore, it was observed in Figure 9 that the internal configurational force at the interface vanishes around the nucleation and only exists where the phase field $\sigma$ is still present.
Electron Microscopy Image Phase field $\varphi(\boldsymbol{x},t)$
@@
In the electron microscopy images presented by You et al. [25], the LNCaP cell line noticeably forms a greater number of nucleations in the cell, and the nucleations appear in higher frequency compared to those in PC3 cells. This can be seen in Figure 13 and is compared to the results from our simulation seen in Figure 8. Thus, we demonstrate we can account for these morphological changes in our simulations by introducing a higher rate of reaction for the degradation of the cyto phase field $\varphi$.
Electron Microscopy Image Phase field $\varphi(\boldsymbol{x},t)$
@@
4.1. Quantitative comparison between electron microscopy images and simulations
To quantitatively compare the scanning electron microscopy images or confocal microscopy images with the simulations exhibiting similar features, we calculate the area. The measurement of area will enable a comparison in the progression of shrinkage of the cell relative to the simulation. The comparison of the area does not indicate similarity in blebbing and fragmentation. For the simulated data, the area is calculated by integrating over the domain at each time point, giving
(19) $A(t)=\frac{1}{|\Omega|}\int_{\Omega}\varphi(\mathbf{x},t)\,\mathrm{d}v.$
For comparing across different length scales for the simulations, we scale the area by the following factor $A/A_{0}$, where $A_{0}=A(t=0)$ (initial condition). We provide plots corresponding to the simulations found in Figures 11 and 13. The volume plot for Figure 12 has not been included, as the same parameters produce the same overall degradation in the volume. Analysing blebs, cavities, and fragments, the initial conditions need to be taken into account, as the difference between the formation of cavities and blebs can be influenced by the initial conditions.
$A/A_{0}$
Time (t~)($\tilde${t})
$A/A_{0}$
Time (t~)($\tilde${t})
Figure 14. (A) Scaled area over time for the simulation using parameter values $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=1.5$, $\ell^{\varphi}_{r}=960$, $\ell^{\sigma}_{r}=0.28$, and $k_{2}=8.0$ (see Table 1 for complete list of parameters). (B) same as (A) with the following parameter values $\ell_{\varphi}=1.333\times 10^{-4}$, $\ell_{2}=5.0$, $\ell^{\varphi}_{r}=4240$, $\ell^{\sigma}_{r}=1.3$, and $k_{2}=1.0$ (see Table 1 for full list of parameter values).
Figure 14 illustrates how the scaled area $(A/A_{0})$ for $\varphi$ (cyto phase) evolves over time (t~)($\tilde${t}) with Figure 14 (A) as the area over time for the simulation presented in Figure 3 and Figure 14 (B) present area over time for the simulation included in Figure 8. These calculations for area can be directly compared to the corresponding experiments to infer parameters in the model these parameters can be identified and assist in understanding therapeutic effectiveness.
@@ 5. Conclusions
Apoptosis proceeds through two well-established signalling routes, that is, the intrinsic, mitochondria-mediated pathway and the extrinsic, receptor-mediated pathway, as discussed by Lossi [18] and Mustafa et al. [20]. Both mechanisms ultimately result in the activation of capases, which coordinate the enzymatic disassembly of the cytoskeleton, organelles, and nucleus. In our formulation, this biochemical activation is represented by a spatially distributed cytotoxic phase, coupled to a cyto phase describing the cell itself. The model treats the activation field in a pathway-agnostic manner, focusing on its mechanical and morphological consequences rather than its molecular origin.
We developed a phase-field framework for biochemical reactions, given by the system in Equations (15), and implemented it in the Dedalus package [4]. Variational derivation yielded a coupled set of reaction–diffusion and configurational-mechanics equations capable of reproducing characteristic apoptotic morphologies and topological transitions. Simulations of the cyto phase $\varphi$ interacting with the cytotoxic phase $\sigma$ reveal finger formation, nucleation, fragmentation, and shrinkage. The parameters $\ell_{\varphi}$, $\ell_{r}^{\varphi}$, $\ell_{r}^{\sigma}$, $k_{2}$, and $\ell_{2}$ control the onset and extent of these topological transitions: decreasing $\tilde{\epsilon}$ suppresses fragmentation, while larger $\ell_{r}^{\varphi}$, $\ell_{r}^{\sigma}$, $k_{2}$, and $\ell_{2}$ enhance cavity and finger formation. Internal configurational-force analysis identifies zones of highest degradation, matching the protrusive regions observed experimentally.
Qualitative comparison with electron microscopy images from You et al. [25], where KML001 induces apoptosis and autophagic cell death in prostate cancer cells via oxidative stress, shows morphological agreement to a great extent. The model captures the same qualitative sequence of surface deformation, cavity formation, and eventual fragmentation observed in PC3 and LNCaP cells, confirming its ability to reproduce experimentally reported apoptotic morphologies. The discrepancies seen between the experimental and simulated data are due to the highly heterogeneous nature of the cytoplasm and the cell as a whole. Furthermore, the simulation utilises a simple initial condition where the cell is assumed to be a circular disk, in reality, cells appear in a myriad of shapes and sizes. Through analysing the area of the cell over time for the microscopy image and the simulation, a method of extracting and comparing measurements has been presented. Further work is required to obtain the area from scanning microscopy images and more biological data for parameter inference. To improve the quantitative comparison, focus should be placed on the fragmentation and blebbing. For analysis, blebbing measurements can be acquired for the width of the blebs and provide a distribution for bleb width over time, and how these compare. The amount of fragmentation can be determined by counting the individual fragments around each cell.
@@ Appendix A Configurational mechanics
The evolution of the interfaces may be described by configurational mechanics. This approach, proposed by Gurtin [13]. Configurational forces are connected to the integrity of the material in a broad sense, explaining the evolution of material and immaterial interfaces. In what follows, we use the configurational balance suggested by Fried [11] and Espath & Calo [9].